OpenStudy (anonymous):

please help me? x^2-10x+25=54 (the answer is x=5(+/-)3√6 )

jimthompson5910 (jim_thompson5910):

Answers are \[\large x=5+3\sqrt{6} \ \ \textrm{or} \ \ x=5-3\sqrt{6}\] See attached image for how I got those answers.

jimthompson5910 (jim_thompson5910):

Note: you can condense the answers above into \[\large x=5\pm3\sqrt{6}\]